• You are given Young's Modulus of Elasticity of 200 GPa and 80 GPa for steel and bronze respectively. Given a cable of steel wire with 4 mm diameter and eight bronze wires of the same size. Find the safe load for the cable and also find the extension under this load on length 43 m. The allowable stress in the bronze wire is 55 MPa. A. 7256.00 N; 29.60 mm
  •  B. 7257.08 N; 29.56 mm
  •  C. 7256.02 N; 30.10 mm
  •  D. 7257.07 N; 30.00 mm
1
Expert's answer
2021-07-07T08:39:56-0400

"F=pS=7256.02~N,"

"E_s+8E_b=\frac F {9S}\frac l{\Delta l},"

"\Delta l=\frac{4Fl}{9\pi d^2(E_s+8E_b)}=30.1~mm."